# M2IST: Multi-Modal Interactive Side-Tuning for Memory-efficient Referring Expression Comprehension

### v1

@@ 4.2 Main Results
In summary, M2IST is Pareto-optimal in terms of accuracy, parameter efficiency, and memory efficiency. By tuning only 3.19M encoder parameters (2.11% of fully fine-tuning) and requiring 15.44GB of GPU memory (39.61% of fully fine-tuning), M2IST makes fine-tuning a strong REC model on a single NVIDIA 3060 GPU (16GB).
4.3 Ablation Study and Analysis
@@
Table 3: Ablation on different components in M3ISA. Without adding any component of M3ISA, it can be viewed as freezing the pre-trained encoder parameters and only training the V-L Encoder.
Effects of Different Components of M3ISA.
Table 3 presents the performance of using different components of M3ISA. We can see that: (1) Freezing the encoders and only training the V-L Encoder leads to much greater performance degradation (Table 3 (a)), indicating a significant domain gap between the pre-trained domains of the two encoders and the REC domain. (2) Fine-tuning single-modality adapters (LEA/VEA) significantly enhances performance compared to using frozen encoders (Table 3 (b,c)). Specifically, VEA provides greater performance improvement compared to LEA, suggesting that adapting visual representation plays a more crucial role in object perception and localization than language representation. (3) Combining LEA and VEA yields similar performance to using IEA alone (Table 3 (d,e)). This indicates that using either can bring around 6% accuracy improvement compared to freezing the encoders. (4) Incorporating LEA, VEA, and IEA into M3ISA results in an average improvement of 8.10% across the three sets of RefCOCO, achieving the best performance among these ablation variants (Table 3 (f)). It is worth noting that fine-tuning each ablation variant of M3ISA incurs at most an additional 1.12GB of GPU memory compared to freezing the encoder, demonstrating the memory efficiency of M2IST (see Section 3.3).
@@
(d) IEA+IEA LEA+VEA 3.19 15.44 81.09 82.29 77.98
Table 4: Effects of different mixing strategies of M3ISA. "VEA+LEA" and "IEA+IEA" refer to adopting the intra-modality adapters and the inter-modality adapters, respectively.
Effects of Different Mixing Strategies of M3ISA.
Table 4 demonstrates the impact of various adapter combination forms (i.e., mixing strategies). The findings are as follows: (1) Transferring pre-trained single-modality knowledge to the REC domain (e.g., LEA+VEA) is more effective in accurately locating the referred object than merely achieving cross-modality interaction (e.g., IEA+IEA) (Table 4 (a,b)). (2) Combining intra-modality adapters and inter-modality adapters enhances performance, indicating that joint transfer of pre-trained single-modality knowledge and cross-modality interaction aids in accurately localizing referred objects by text descriptions (Table 4 (a,b,c,d)). This observation aligns with findings from other challenging vision-language tasks Xu et al. (2023); Zhou and Long (2023b), suggesting that combining deep inter-modality fusion with intra-modality adaptation improves performance. (3) The best performance among the M3ISA variants is achieved by first connecting the vision and language encoders with IEAs, and then adapting the interacted features and single-modality features to the REC domain with VEA and LEA (Table 4 (a,b,c,d)).
Effects of Different Insertion Forms of M3ISA.
As depicted in Figure 3 and Table 5, we evaluate the impact of integrating M3ISAs with different insertion forms on performance and GPU memory usage. (1) Side insertion yields the best performance. We suppose that implementing M3ISAs on side networks enhances the alignment between the referring sentence and the referred object, resulting in improved localization performance. (2) All three insertion forms reduce GPU memory usage to varying degrees. Incorporating M3ISAs into the side networks consumes the least amount of GPU memory. This is because the gradients backpropagate through the lightweight M3ISAs instead of heavy encoders. This aligns with the memory efficiency advantage mentioned in Section 3.3.
@@
Table 5: Effects of different insertion forms of M3ISA. "Sequential" and "Parallel", and "Side" correspond to (a), (b), and (c) in Figure 3, respectively.
4.4 Qualitative Results
To investigate the impact of cross-modality interaction facilitated by M3ISAs, we visualize the attention maps from the V-L Encoder. We compare M3ISA with its variant presented in Table 4 (a) under various scenarios, shown in Figure 4. It is evident that M3ISA can handle diverse REC cases, indicating that the enhanced cross-modality interaction enabled by the IEA allows for effective comprehension of complex semantic information.
5 Conclusion
In this paper, we present Multi-Modal Interactive Side-Tuning (M2IST), a parameter- and memory-efficient tuning method for REC. We introduce Mixture of Multi-Modal Interactive Side Adapters (M3ISA) to efficiently transfer pre-trained single-modality knowledge and facilitate cross-modality interaction between vision and language encoders. During fine-tuning, we freeze the pre-trained vision and language encoders and update M3ISAs on side networks, achieving efficient tuning for REC. By updating only 3.14M encoder parameters (2.11% of full fine-tuning) and using 15.44GB of GPU memory (39.61% of full fine-tuning), M2IST achieves competitive performance compared to full fine-tuning methods and outperforms other PETL methods across three benchmarks.
6 Limitations
In this work, we implement our M2IST on the mainstream transformer-based architecture for referring expression comprehension, comprising a pre-trained Vision Encoder and Language Encoder. With the rapid development of multi-modal large language models (MLLMs), applying M2IST to MLLMs (e.g., LLaVA Liu et al. (2023) and InstructBLIP Dai et al. (2023)) could potentially further enhance their reasoning capabilities in complex scenarios Pan et al. (2024a, b). Due to limited computational resources, our experiments were conducted only using ResNet-50 with the DETR encoder. Future work will involve more extensive experiments with ViT-L Dosovitskiy et al. (2021) and Swin-L Liu et al. (2021) backbones to fully explore the scalability and potential of M2IST.
Appendix A Details of REC Datasets
To verify the effectiveness and efficiency of our method, we conduct experiments on the following REC benchmarks as follows:
  • RefCOCO Yu et al. (2016) consists of 19,994 images with 142,210 referring expressions for 50,000 objects. The RefCOCO dataset is officially split into train, validation, testA, and testB sets containing 120,624, 10,834, 5,657, and 5,095 expressions, respectively.
  • RefCOCO+ Yu et al. (2016) includes 19,922 images with 141,564 referring expressions for 49,856 objects. Compared to RefCOCO, the referring expressions in RefCOCO+ focus more on attributes of the referred objects, such as color and shape, without including any positional words.
  • RefCOCOg Mao et al. (2016); Nagaraja et al. (2016) contains 25,799 images with 95,010 referring expressions for 49,822 objects. Compared to RefCOCO and RefCOCO+, the referring expressions in RefCOCOg are typically longer, averaging almost twice the length of those in the other two datasets. RefCOCOg has two commonly used split strategies: the google split Mao et al. (2016) (-g) and the umd split Nagaraja et al. (2016) (-u). Following previous work Deng et al. (2021); Yang et al. (2022); Zhu et al. (2022), we conduct experiments on both RefCOCOg-g (val-g) and RefCOCOg-u (val-u and test-u).
Appendix B More Implementation Details
Model Weights.
The Vision Encoder is initialized with the backbone (i.e., ResNet-50 He et al. (2016)) and encoder weights from DETR Carion et al. (2020), which is pre-trained on the MS-COCO dataset Lin et al. (2014). Specifically, during the pre-training of the Vision Encoder, images from the validation and test sets of RefCOCO/+/g that overlap with MS-COCO Lin et al. (2014) are excluded. The Language Encoder is initialized with BERT-base Devlin et al. (2018), pre-trained on the BookCorpus Zhu et al. (2015) and English Wikipedia Devlin et al. (2018). The Vision-Language (V-L) Encoder is initialized using Xavier initialization. The proposed M3ISAs are initialized with Kaiming normal initialization.
Hyper-parameters Settings.
M3ISAs are inserted into the transformer encoder layers at the same indices as those in the Vision Encoder and Language Encoder, and relevant ablation study is conducted in Table 6. The bottleneck dimensions of the Visual Embedding Adapter (VEA) and Language Embedding Adapter (LEA) are set to 128, while the interaction dimension $C_{i}$ of the Interaction Embedding Adapter (IEA) is 256. A relevant ablation study on these hyperparameters is presented in Table 7. The scaling factor $s$ for all adapters is set to 0.1.
Training Details.
For RefCOCO Yu et al. (2016) and RefCOCOg Mao et al. (2016); Nagaraja et al. (2016) datasets, the entire network is trained for 90 epochs using the AdamW optimizer Loshchilov and Hutter (2019), with a learning rate of $10^{-4}$ for the V-L Encoder and $10^{-5}$ for the M3ISAs. The weight decay is $10^{-4}$, and the learning rate is reduced by a factor of 10 after 60 epochs. While for RefCOCO+ Yu et al. (2016) dataset, the network is trained for 180 epochs with the same learning rates and weight decay, but the learning rate is decreased by a factor of 10 after 120 epochs. We conduct all experiments on one A800 GPU.
Appendix C Training Objective
Following most transformer-based REC methods Deng et al. (2021); Yang et al. (2022); Su et al. (2023), the training loss function is a combination of the widely used smooth L1 loss and GIoU loss. Specifically, the prediction is donated as $\mathbf{b}=(x,y,w,h)$, and the normalized ground-truth box as $\hat{\mathbf{b}}=(\hat{x},\hat{y},\hat{w},\hat{h})$. The training objective is:
Appendix D Details of Baseline PETL Methods
This section furnishes additional details of the PETL baselines employed in our primary manuscript. Notably, all these baselines follow the same base architecture, wherein the Vision Encoder and Language Encoder remain fixed, while the V-L Encoder and the newly added parameters are updated during fine-tuning.
  • Adapter Houlsby et al. (2019): We incorporate standard adapters behind the Multi-head Attention (MHA) layers and Feed-Forward Networks (FFN) in both Vision Encoder and Language Encoder. Consistent with our M3ISAs, we set the bottleneck dimensions of these adapters to 128.
  • LoRA Hu et al. (2022): We incorporate trainable matrices in parallel to the weight matrices in MHA and FFN in both Vision Encoder and Language Encoder. Consistent with our M3ISAs for a fair comparison, we employ a LoRA rank of $r=128$ for both vision and language branch.
  • AdaptFormer Chen et al. (2022): We add adapters in parallel to MHA and FFN in both Vision Encoder and Language Encoder. Similar to Adapter Houlsby et al. (2019), we set bottleneck dimensions of AdaptFormer to 128 for both vision and language branch.
  • CM Adapter Jiang et al. (2022): We sequentially insert CM Adapters after the MHA and FFN layers of the encoder layers with the same indices as in Vision Encoder and Language Encoder. Consistent with our M3ISAs, we set the bottleneck dimensions of CM Adapter to 128, and the weight-sharing dimensions of CM Adapter to 256.
  • MRS-Adapter Yuan et al. (2023): We add MRS-Adapters in parallel to FFN in both Vision Encoder and Language Encoder, according to their basic designs. Similar to CM Adapter Jiang et al. (2022), we set the bottleneck dimensions of MRS-Adapter to 128, and the weight-sharing dimensions of MRS-Adapter to 256.
Appendix E Additional Ablation Study
In this section, we conduct more ablative experiments to further explore the impact of various factors in M2IST. All experiments are performed on three sets of RefCOCO Yu et al. (2016) dataset.
Effects of Different Insertion Positions of M3ISA.
As illustrated in Table 6, we further investigate the impact of introducing M3ISAs at different positions within the pre-trained Vision Encoder and Language Encoder. The Vision Encoder and Language Encoder consist of 6 and 12 transformer encoder layers, respectively, and the IEA needs to be inserted into the encoder layers at the same indices. We explore three common insertion forms, as shown in Table 6 (a-c). It is evident that inserting M3ISAs in parallel to the deeper encoder layers of the pre-trained Language Encoder results in better performance. We suggest that deeper encoder layers contain richer semantic features, and establishing cross-modality interaction on this basis helps the model learn finer region-text alignment, thereby achieving better localization performance.
@@ Appendix E Additional Ablation Study
Table 6: Effects of Different Insertion Positions of M3ISA. The Vision Encoder and Language Encoder consist of 6 and 12 transformer encoder layers, respectively. "$1\rightarrow 6$" denotes the addition of M3ISAs in the 1st through 6th transformer encoder layers.
Effects of Different Hyper-parameter Settings of M3ISA.
We first ablate the bottleneck dimensions $C_{d}$ of the intra-modality adapters (see Table 7 (a,b,c)), and follow the design shown in Table 4 (a). $C_{d}$ determine the number of tunable parameters introduced by M3ISA. As shown in Table 7, higher $C_{d}$ introduces more parameters, and the performance consistently increases when $C_{d}$ increases up to 128. Thus, we select the $C_{d}$ as 128. We further ablate the impact of changing the interaction dimensions $C_{i}$ of inter-modality adapters (i.e., IEA), and follow the paradigm of Table 4 (d). As depicted in Table 7 (e, f, g), deeper cross-modality interaction results in an increase in tunable parameters and performance. Thus, $C_{d}$ and $C_{i}$ are set to 128 and 256, respectively, to achieve the optimal trade-off among accuracy, number of tunable parameters, and GPU memory consumption. It is worth noting that all ablative variants exhibit a remarkable level of memory efficiency, as they consume less than 16GB of GPU memory. This observation is consistent with the memory efficiency advantage highlighted in Section 3.3.
# $C_{d}$/$C_{i}$ Params.$\downarrow$ Mem.$\downarrow$ RefCOCO
@@
(a) 32 0.85 15.53 77.46 77.91 73.96
(b) 64 1.64 15.53 79.37 80.13 75.96
(c) 128 3.21 15.64 80.69 81.76 76.43
(e) 64 2.00 15.34 77.31 77.87 73.27
(f) 128 2.40 15.35 79.26 79.58 74.60
(g) 256 3.19 15.44 81.35 82.29 77.98
Table 7: Effects of different hyper-parameter settings of M3ISA. "$C_{d}$" denotes the bottleneck dimensions of VEA and LEA, while "$C_{i}$" represents the interaction dimensions of IEA. In (a)-(c), we simply use intra-modality adapters (see Table 4 (a)) to find the most suitable $C_{d}$. Subsequently, in (d)-(f), we keep the $C_{d}$ obtained from (a)-(c) fixed and explore different values for $C_{i}$ to identify the most appropriate $C_{i}$.
Appendix F More Visualization Results
In this section, we present more visualization of the attention maps from V-L Encoder under different mixing strategies (i.e., without interaction and with interaction). As depicted in Figure 5, the interaction between the vision and language encoder, facilitated by M3ISAs, allows the model to focus more effectively on the referred objects in diverse referring expression comprehension (REC) cases, including object appearance attributes, human actions, and spatial relations.
Figure 5: More visualizations of attention maps from V-L Encoder with different mixing strategies of M3ISA. Cases include object appearance attributes (blue words), human actions (green words), and spatial relations (red words).
Appendix G Pseudocode of M2IST
We present the PyTorch-like pseudocode of our proposed M2IST in Algorithm 1 to help to better understand the whole process.
Algorithm 1 PyTorch-like pseudocode of M3ISAs in vision and language encoder layers.
# Frozen the pre-trained encoders except all adapters
for name, p in model.named_parameters():
if "adapter" in name:
p.requires_grad = True
else:
p.requires_grad = False
# Define the VEA and LEA Module, taking VEA for example.
class VEA(nn.Module):
def __init__(self, d_model, bottleneck, dropout, adapter_scalar):
super().__init__()
self.n_embd = d_model
self.down_size = bottleneck
self.down_proj = nn.Linear(self.n_embd, self.down_size)
self.non_linear_func = nn.ReLU()
self.visual_up_proj = nn.Linear(self.down_size, self.n_embd)
self.dropout = dropout
self.scale = adapter_scalar
def forward(self, x):
down = self.down_proj(x)
down = self.non_linear_func(down)
down = nn.functional.dropout(down)
output = up * self.scale
return output
# Define the IEA Module.
class IEA(nn.Module):
def __init__(self, vis_d_model, text_d_model, bottleneck, share_bottleneck, share_up, adapter_scalar):
super().__init__()
self.vis_d_model = vis_d_model
self.text_d_model = text_d_model
self.up_size = bottleneck
self.share_size = share_bottleneck
self.share_up = share_up
self.scale = adapter_scalar
self.text_down_proj = nn.Linear(self.text_d_model, self.share_size)
self.vis_down_proj = nn.Linear(self.vis_d_model, self.share_size)
self.up_proj_share = nn.Linear(self.share_size, self.share_up)
self.text_up_proj = nn.Linear(self.share_size, self.text_d_model)
self.vis_up_proj = nn.Linear(self.share_size, self.vis_d_model)
def forward(self, text_x, vis_x):
vis_down = self.vis_down_proj(vis_x)
text_down = self.text_down_proj(text_x)
text_up = self.up_proj_share(text_x)
vis_up = self.up_proj_share(vis_x)
return up
IEA_out = []
LEA_out = []
VEA_out = []
# Multi-Modal Interactive Side-Tuning
for i in range(layers):
IEA_out = LEA(text_mha_output,vis_mha_output)
LEA_out = LEA(text_ffn_out)
VEA_out = VEA(vis_ffn_out)
IEA_out.append(IEA_out)
LEA_out.append(LEA_out)
VEA_out.append(VEA_out)
final_text_feature = text_feature + sum(IEA_out) + sum(LEA_out)
final_vis_feature = vis_feature + sum(IEA_out) + sum(VEA_out)

### v2

@@ IV-C Main Results
In summary, M2IST is Pareto-optimal in terms of accuracy, parameter efficiency, memory efficiency, and time efficiency. By tuning only 3.19M encoder parameters (2.11% of fully fine-tuning) and requiring 15.44GB of GPU memory (39.61% of fully fine-tuning) and 63.46% of the full fine-tuning time, M2IST makes fine-tuning a strong REC model on a single NVIDIA 3060 GPU (16GB).
Comparison on Phrase Grounding Task
To demonstrate the generalizability of our M2IST, we have conducted experiments with M2IST on the task of phrase grounding.
Phrase grounding is a challenging vision-language task that given a noun phrase, output the corresponding single or multiple object detection bounding boxes. If an entire sentence is input, then it’s about localizing all the noun phrases included in the sentence. Here, we use the same base architecture in our paper for phrase grounding on ReferItGame dataset [57]. Table IV demonstrates the generalization ability of our M2IST in understanding multi-object scenarios, while also exhibiting significant parameter and memory efficiency during fine-tuning.
IV-D Ablation Study and Analysis
In this sub-section, we investigate the impact of various factors in M2IST. All experiments in this section are performed on three sets of RefCOCO [22] dataset.
Effects of Different Components of M3ISA
We present the efficiency and performance of various components of M3ISA to examine their effects, as shown in Table V.
We can see that: (1) Freezing the encoders and only training the V-L Encoder leads to much greater performance degradation (Table V (a)), indicating a significant domain gap between the pre-trained domains of the two encoders and the REC domain. (2) Fine-tuning single-modality adapters (LEA/VEA) significantly enhances performance compared to using frozen encoders (Table V (b,c)). Specifically, VEA provides greater performance improvement compared to LEA, suggesting that adapting visual representation plays a more crucial role in object perception and localization than language representation. (3) Combining LEA and VEA yields similar performance to using IEA alone (Table V (d,e)). This indicates that using either can bring around 6% accuracy improvement compared to freezing the encoders. (4) Incorporating LEA, VEA, and IEA into M3ISA results in an average improvement of 8.10% across the three sets of RefCOCO, achieving the best performance among these ablation variants (Table V (f)). It is worth noting that fine-tuning each ablation variant of M3ISA incurs at most an additional 1.12GB of GPU memory compared to freezing the encoder, demonstrating the memory efficiency of M2IST (see Section III-C).
TABLE V: Ablation on different components in M3ISA. Without adding any component of M3ISA, it can be viewed as freezing the pre-trained encoder and only training the V-L Encoder.
@@ IV-D Ablation Study and Analysis
TABLE VI: Effects of different mixing strategies of M3ISA. "VEA+LEA" and "IEA+IEA" refer to adopting the intra-modality adapters and the inter-modality adapters, respectively.
@@
(d) IEA+IEA LEA+VEA 3.19 15.44 81.35 82.29 77.98
Effects of Different Mixing Strategies of M3ISA
In Table VI, to further investigate the effects of different adapter combination forms (i.e., mixing strategies), we present the performance of adopting intra-modality adapters or inter-modality adapters in parallel with different pre-trained layers (MHA and FFN). The findings are as follows: (1) Transferring pre-trained single-modality knowledge to the REC domain (e.g., LEA+VEA) is more effective in accurately locating the referred object than merely achieving cross-modality interaction (e.g., IEA+IEA) (Table VI (a,b)). (2) Combining intra-modality adapters and inter-modality adapters enhances performance, indicating that joint transfer of pre-trained single-modality knowledge and cross-modality interaction aids in accurately localizing referred objects by text descriptions (Table VI (a,b,c,d)). This observation aligns with findings from other challenging vision-language tasks [16, 8], suggesting that combining deep inter-modality fusion with intra-modality adaptation improves performance. (3) The best performance among the M3ISA variants is achieved by first connecting the vision and language encoders with IEAs, and then adapting the interacted features and single-modality features to the REC domain with VEA and LEA (Table VI (a,b,c,d)).
Effects of Different Insertion Forms of M3ISA
As depicted in Figure 3 and Table VII, we evaluate the impact of integrating M3ISAs with different insertion forms on performance and GPU memory usage. From Table VII, we can observe that: (1) Side insertion yields the best performance. We suppose that implementing M3ISAs on side networks enhances the alignment between the referring sentence and the referred object, resulting in improved localization performance. (2) In terms of fine-tuning efficiency, all three insertion forms contribute to a reduction in GPU memory usage to varying degrees. It is evident that incorporating M3ISAs into the side networks consumes the least amount of GPU memory. This is because the gradients backpropagate through the lightweight M3ISAs instead of heavy encoders. This aligns with the memory efficiency advantage mentioned in Section III-C.
TABLE VII: Effects of different insertion forms of M3ISA. "Sequential" and "Parallel", and "Side" correspond to (a), (b), and (c) in Figure 3, respectively.
@@
TABLE VIII: Effects of Different Insertion Positions of M3ISA. The Vision Encoder and Language Encoder consist of 6 and 12 transformer encoder layers, respectively. "$1\rightarrow 6$" denotes the addition of M3ISAs in the 1st through 6th encoder layers.
@@
Effects of Different Insertion Positions of M3ISA
As illustrated in Table VIII, we further investigate the impact of introducing M3ISAs at different positions within the pre-trained Vision Encoder and Language Encoder. The Vision Encoder and Language Encoder consist of 6 and 12 transformer encoder layers, respectively, and the IEA needs to be inserted into the encoder layers at the same indices. We explore three common insertion forms, as shown in Table VIII (a-c). It is evident that inserting M3ISAs in parallel to the deeper encoder layers of the pre-trained Language Encoder results in better performance. We suggest that deeper encoder layers contain richer semantic features, and establishing cross-modality interaction on this basis helps the model learn finer region-text alignment, thereby achieving better localization performance.
TABLE IX: Effects of different interaction dimensions of M3ISA. "$C_{i}$" represents the interaction dimension of Interaction Expert Adapter (IEA) in M3ISAs.
# $C_{i}$ Params.$\downarrow$ Mem.$\downarrow$ RefCOCO
@@
(a) 64 2.00 15.34 77.31 77.87 73.27
(b) 128 2.40 15.35 79.26 79.58 74.60
(c) 256 3.19 15.44 81.35 82.29 77.98
Effects of Different Interaction Dimensions of M3ISA
We further ablate the impact of changing the interaction dimensions $C_{i}$ of inter-modality adapters (i.e., IEA), and follow the paradigm of Table VI (d). As depicted in Table IX, deeper cross-modality interaction provides better vision-language channel-level alignment, thus resulting in an increase in performance. Thus, $C_{i}$ is set to 256 to achieve the optimal trade-off among accuracy, number of tunable parameters, and GPU memory consumption. It is worth noting that all ablative variants exhibit a remarkable level of memory efficiency, as they consume less than 16GB of GPU memory. This observation is consistent with the memory efficiency advantage highlighted in Section III-C.
TABLE X: Effects of combining M2IST with LoRA. "Full" represents fully fine-tuning the baseline modal.
Methods Params.$\downarrow$ Mem.$\downarrow$ RefCOCO
(M) (GB) val testA testB
Full 151 38.95 80.32 82.67 75.24
LoRA 2.37 20.37 77.57 78.22 73.37
M2IST 3.19 15.44 81.35 82.29 77.98
M2IST+LoRA 6.46 21.68 81.83 82.83 77.54
Effects of Combining M2IST with LoRA
To evaluate the extensibility of our M2IST, we combine it with the widely used PETL method, LoRA [13].
As shown in Table X, using LoRA alone leads to performance degradation. This is because LoRA integrates two pairs of tunable low-rank decomposed weight matrices into each encoder layer of the Vision Encoder and Language Encoder separately, without any multi-modality interaction during the feature extraction stage. Combining M2IST with LoRA performs comparably to or even better than using M2IST alone, demonstrating the extensibility of our approach. However, as discussed in Section III-C, such a combination may significantly increase the number of updated parameters and GPU memory usage, thus undermining the memory efficiency advantage of M2IST. Therefore, our M2IST achieves the optimal performance and efficiency trade-off in this comparison.
IV-E Qualitative Results
To investigate the impact of cross-modality interaction facilitated by M3ISA, we visualize the attention maps from the V-L Encoder. We compare M3ISA (referred to as "With Interaction") with its variant presented in Table VI (referred to as "Without Interaction") across various scenarios to assess their effectiveness in understanding challenging cases, such as human action recognition and spatial relation reasoning, as shown in Figure 4.
It is clear that M3ISA effectively addresses diverse REC cases, particularly in spatial relations. For instance, in the sixth case, the "With Interaction" approach successfully directs attention to the front region associated with the referring sentence "front pizza". This is especially evident in the last case, where "With Interaction" focuses more on the area corresponding to the referring sentence "right person in air", while "Without Interaction" fails to identify the correct region. These visualization results demonstrate that the enhanced cross-modality interaction facilitated by the IEA enables effective comprehension of complex semantic information.
V Conclusion
In this paper, we introduce Multi-Modal Interactive Side-Tuning (M2IST), an efficient tuning method designed for referring expression comprehension. Based on this framework, we introduce Mixture of Multi-Modal Interactive Side Adapters (M3ISA) to efficiently transfer pre-trained single-modality knowledge and facilitate cross-modality interaction between vision and language encoders. During fine-tuning, we freeze the pre-trained vision-language foundation models and update M3ISAs on side networks, achieving efficient tuning for REC. By updating only 3.14M encoder parameters (2.11% of full fine-tuning) and using 15.44GB of GPU memory (39.61% of full fine-tuning), M2IST achieves competitive performance compared to full fine-tuning methods and outperforms other PETL methods across three benchmarks.